### v1

@@ 3.2 Circuit Architecture Search for Data Embedding
Next, we check if our scheme can be exploited to find better embedding quantum circuits. For classical data tasks, embedding quantum circuits play a crucial role in the performance, since the data distribution in the Hilbert space is determined by them. Previous works introduced techniques to increase the expressivity in terms of Fourier analysis schuld2021effect ; shin2023exponential . Still, there is room for investigation in guidelines for constructing performant embedding quantum circuits. Hence, we here propose a new embedding strategy based on our light-cone feature selection technique.
@@
For the proof-of-concept demonstration of this scheme, we again work on the parity-based binary classification task. In the numerical experiments, we consider the dataset with $A_{p}=\{1,5\}$ denoted as parity(1,5). Table 1 shows the KTA values for the cases with and without the re-ordering technique. Here, the optimization of local kernels are done three times for both cases. It turns out that the KTAs increase from -0.033 to 0.987 (-0.015 to 0.560) for QFKs (PQKs). The performance improvement can be interpreted in terms of feature selection. Figure 3 shows the importance scores for these cases. When the technique is not employed, every feature is evenly regarded as important for QFKs and PQKs. Together with the KTA values, this indicates that the optimization works poorly because the restriction on entangling gates limits the spread of light-cones. Note that a few local terms in PQKs can cover features $x_{1}$ and $x_{5}$, but they are so far away that optimization could not lead to good prediction. On the other hand, QFKs and PQKs can successfully select the relevant features, suggesting that the re-ordering facilitates the optimization and results in the better performance.
@@
Moreover, we work on the benchmark dataset, the breast-cancer dataset, to see the performance for a more practical task. This is a classical dataset for binary classification where 212 data points are labeled as “malignant” and 357 as “benign”. In the numerical experiments, we choose $d$($=12$) features out of 30 and then rescale the data so that it has a mean zero and unit variance. The distribution of each feature is shown in Figure 4 (a), in which we show the L1 norm of “malignant” and “benign” distributions $\delta$ to show the discrepancy (i.e., separability) of them. The dataset is taken from scikit-learn scikit-learn . Also, we employ the same quantum circuits as the parity-based classification tasks.
@@ 3.3 Compression of Quantum Machine Learning Models
Another application where our proposal can be utilized is the compression of quantum circuit models. The model compression aims to reduce the size of machine learning models, e.g., the number of parameters in neural networks cheng2017survey ; as a result, resource efficiency, faster inference and better generalization can be realized.
Especially, pruning is a powerful tool to reduce the model size in classical machine learning. With the pruning technique, one can construct a compressed model that requires less resources to execute inference and is comparable to the original model in the performance. For example, parameters of neural networks are pruned based on their magnitudes frankle2018lottery (i.e., the parameters with small values are removed). Indeed, the magnitude-based pruning method has been explored in variational quantum algorithms sim2021adaptive ; wang2022quantumnas . On the other hand, careful consideration is needed when applying the magnitude-based pruning to data re-uploading quantum circuits. More concretely, this technique fails to prune the embedding layer, suggesting its inability to reduce the model width (i.e., the number of qubits). This also means that the reduction in the depth might not be huge as the embedding layers will never be removed.
@@ 3.4 Subspace Feature Selection for Quantum Data
To see the effectiveness of our proposal, we consider a toy binary classification task with a synthetic quantum dataset $\{\rho_{i},y_{i}\}$. Here, the input quantum state is represented as $\rho_{i}=U(\bm{\theta}^{*})\rho_{0,i}U^{\dagger}(\bm{\theta}^{*})$, where $\rho_{0,i}=\rho_{0,i}^{(1)}\otimes\rho_{0,i}^{(2)}\otimes\ldots\otimes\rho_{0,i}^{(n)}$ is the tensor product of single-qubit Haar random states $\{\rho_{0,i}^{(l)}\}_{l=1,\ldots,n}$ and $U(\bm{\theta}^{*})$ denotes the unitary operator with certain fixed parameters $\bm{\theta}^{*}$. Then, the label is determined by $y_{i}=\text{sign}(\mathrm{Tr}[Z_{m}\rho_{0,i}])$ with the Pauli $Z$ operator acting on the $m$-th qubit $Z_{m}$; see Figure 6 (a) for the details of this numerical experiment. We construct the dataset inspired by actual quantum data tasks; like the quantum phase recognition tasks in Ref. lake2022exact , we assume that there exist a unitary operator that can transform input quantum state $\rho_{i}$ to another state in the form of $\tilde{\rho}_{i,r}\otimes\rho_{irr}$ with a relevant quantum state $\tilde{\rho}_{i,r}$ and an irrelevant state $\rho_{irr}$. By a relevant quantum state, we mean that a simple measurement on that state is enough to identify the label or the property of the original state. In this work, we focus on PQKs using alternating layered ansatzes cerezo2021cost with the number of qubit $n=8,12$ and the circuit depth $L=1,2,3$. Note that we do not take into account QFKs since the kernels are not defined via input quantum states $\rho_{i}$, but via the unitary representation of data $U(\bm{x_{i}},\bm{\theta})$. Also, we set $m=n/2+1$.
@@ 4 Conclusion & Discussion
As described in Section 2, our framework relies on the training of parameters $(\bm{\theta},\bm{\lambda})$ in quantum kernels; this means that the classical computational cost for the quantum kernel estimation scales at least quadratically with respect to the number of training data points. Hence, there is room for investigation in improving its scalability. As the remedy to the similar issue in quantum kernel methods, previous works used techniques such as classical shadow and surrogate models based on random Fourier features huang2021power ; nakaji2022deterministic ; shin2023analyzing . Thus, it would be interesting to explore the improvement of its scalability in this direction.
According to Ref. cerezo2023does , the absence of trainability issues such as barren plateaus and exponential concentrations is strongly tied with classical simulability. This indicates that, for classical data tasks, our technique might not work in the classically-hard regime; even if it works, the compressed model obtained through our method could be classically simulable. On the other hand, this also means that our method can be used as a practical test to see if the QML tasks really requires quantumness. More significantly, our proposal is also applicable to quantum data tasks such as quantum phase recognition cong2019quantum ; lake2022exact ; wu2023quantum and entanglement detection guhne2009entanglement (i.e, quantum physics-related learning tasks). Therefore, our proposal could also be used for constructing efficient settings of such quantum tasks and the investigation of practical quantum advantages.
Acknowledgements.

### v2

@@ 3.2 Circuit Architecture Search for Data Embedding
Next, we check if our scheme can be exploited to find better embedding quantum circuits. For classical data tasks, embedding quantum circuits play a crucial role in the performance, since the data distribution in the Hilbert space is determined by them. Previous works introduced techniques to increase the expressivity in terms of Fourier analysis [52, 53]. Still, there is room for investigation in guidelines for constructing performant embedding quantum circuits. Hence, we here propose a new embedding strategy based on our light-cone feature selection technique.
@@
For the proof-of-concept demonstration of this scheme, we again work on the parity-based binary classification task. In the numerical experiments, we consider the dataset with $A_{p}=\{1,5\}$ denoted as parity(1,5). Table 1 shows the KTA values for the cases with and without the re-ordering technique. Here, the optimization of local kernels are done three times for both cases. It turns out that the KTAs increase from -0.033 to 0.987 (-0.015 to 0.560) for QFKs (PQKs). The performance improvement can be interpreted in terms of feature selection. Figure 4 shows the importance scores for these cases. When the technique is not employed, every feature is evenly regarded as important for QFKs and PQKs. Together with the KTA values, this indicates that the optimization works poorly because the restriction on entangling gates limits the spread of light-cones. Note that a few local terms in PQKs can cover features $x_{1}$ and $x_{5}$, but they are so far away that optimization could not lead to good prediction. On the other hand, QFKs and PQKs can successfully select the relevant features after the re-ordering, suggesting that this technique facilitates the optimization and results in the better performance.
@@
Moreover, we work on the benchmark dataset, the breast-cancer dataset, to see the performance for a more practical task. This is a classical dataset for binary classification where 212 data points are labeled as “malignant" and 357 as “benign". In the numerical experiments, we choose $d$($=12$) features out of 30 and then rescale the data so that it has a mean zero and unit variance. The distribution of each feature is shown in Figure 5 (a), in which we show the L1 norm of “malignant" and “benign" distributions $\delta$ to show the discrepancy (i.e., separability) of them. The dataset is taken from scikit-learn [54]. Also, we employ the same quantum circuits as the parity-based classification tasks.
@@ 3.3 Compression of Quantum Machine Learning Models
Another application where our proposal can be utilized is the compression of quantum circuit models. The model compression aims to reduce the size of machine learning models, e.g., the number of parameters in neural networks [55]; as a result, resource efficiency, faster inference and better generalization can be realized.
Especially, pruning is a powerful tool to reduce the model size in classical machine learning. With the pruning technique, one can construct a compressed model that requires less resources to execute inference and is comparable to the original model in the performance. For example, parameters of neural networks are pruned based on their magnitudes [56] (i.e., the parameters with small values are removed). Indeed, the magnitude-based pruning method has been explored in variational quantum algorithms [57, 58]. On the other hand, careful consideration is needed when applying the magnitude-based pruning to data re-uploading quantum circuits. More concretely, this technique fails to prune the embedding layer, suggesting its inability to reduce the model width (i.e., the number of qubits). This also means that the reduction in the depth might not be huge as the embedding layers will never be removed.
@@ 3.4 Subspace Feature Selection for Quantum Data
To see the effectiveness of our proposal, we consider a toy binary classification task with a synthetic quantum dataset $\{\rho_{i},y_{i}\}$. Here, the input quantum state is represented as $\rho_{i}=U(\bm{\theta}^{*})\rho_{0,i}U^{\dagger}(\bm{\theta}^{*})$, where $\rho_{0,i}=\rho_{0,i}^{(1)}\otimes\rho_{0,i}^{(2)}\otimes\ldots\otimes\rho_{0,i}^{(n)}$ is the tensor product of single-qubit Haar random states $\{\rho_{0,i}^{(l)}\}_{l=1,\ldots,n}$ and $U(\bm{\theta}^{*})$ denotes the unitary operator with certain fixed parameters $\bm{\theta}^{*}$. Then, the label is determined by $y_{i}=\text{sign}(\mathrm{Tr}[Z_{m}\rho_{0,i}])$ with the Pauli $Z$ operator acting on the $m$-th qubit $Z_{m}$; see Figure 7 (a) for the details of this numerical experiment. We construct the dataset inspired by actual quantum data tasks; like the quantum phase recognition tasks in Ref. [59], we assume that there exist a unitary operator that can transform input quantum state $\rho_{i}$ to another state in the form of $\tilde{\rho}_{i,r}\otimes\rho_{irr}$ with a relevant quantum state $\tilde{\rho}_{i,r}$ and an irrelevant state $\rho_{irr}$. By a relevant quantum state, we mean that a simple measurement on that state is enough to identify the label or the property of the original state. In this work, we focus on PQKs using alternating layered ansatzes [19] with the number of qubit $n=8,12$ and the circuit depth $L=1,2,3$. Note that we do not take into account QFKs since the kernels are not defined via input quantum states $\rho_{i}$, but via the unitary representation of data $U(\bm{x_{i}},\bm{\theta})$. Also, we set $m=n/2+1$.
@@ 4 Conclusion & Discussion
As described in Section 2, our framework relies on the training of parameters $(\bm{\theta},\bm{\lambda})$ in quantum kernels; this means that the classical computational cost for the quantum kernel estimation scales at least quadratically with respect to the number of training data points. Hence, there is room for investigation in improving its scalability. As the remedy to the similar issue in quantum kernel methods, previous works used techniques such as classical shadow and surrogate models based on random Fourier features [23, 60, 61]. Thus, it would be interesting to explore the improvement of its scalability in this direction.
According to Ref. [37], the absence of trainability issues such as barren plateaus and exponential concentrations is strongly tied with classical simulability. This indicates that, for classical data tasks, our technique might not work in the classically-hard regime; even if it works, the compressed model obtained through our method could be classically simulable. On the other hand, this also means that our method can be used as a practical test to see if the QML tasks really require quantumness. More significantly, our proposal is also applicable to quantum data tasks such as quantum phase recognition [62, 59, 63] and entanglement detection [64] (i.e, quantum physics-related learning tasks). Therefore, our proposal could also be used for constructing efficient settings of such quantum tasks and the investigation of practical quantum advantages.
Acknowledgements